OpenStudy (juliannamartineez):

what is the solution of the following system? [-2x=y=1 [-4x - 2y = -1

jhonyy9 (jhonyy9):

Julianna please correct the first equation bc. not can having there two equal signs

OpenStudy (juliannamartineez):

-2x - y = 1

jhonyy9 (jhonyy9):

ok -2x -y =1 multiply the first equation by -2 and assume term by term -4x -2y = -1 ---------------- 4x +2y = -2 -4x -2y = -1 ---------------- 0 0 = -3 0 = -3 do you know what mean this result ?

OpenStudy (juliannamartineez):

im confused

jhonyy9 (jhonyy9):

so this mean that this system have not solutions hope helped